# Minimal model program for log canonical pairs on complex analytic spaces

### v1

@@ 3.5. From existence of weak lc model to termination
that defines a sequence of steps of a $(K_{\widetilde{X}}+\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$-MMP over $Y$ around $W$ terminating with a log minimal model $(\widetilde{X}^{\prime\prime\prime},\widetilde{\Delta}^{\prime\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime\prime})$ over $Y$ around $W$. By (iii) in Step 3, $(\widetilde{X}^{\prime\prime\prime},\widetilde{\Delta}^{\prime\prime\prime})$ is a weak lc model of $(\widetilde{X},\widetilde{\Delta})$ over $Y$ around $W$. Since $0\leq\lambda_{n}<\lambda^{\prime}$, we see that $(\widetilde{X}^{\prime\prime\prime},\widetilde{\Delta}^{\prime\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime\prime})$ and $(\widetilde{X}^{\prime\prime\prime},\widetilde{\Delta}^{\prime\prime\prime}+\lambda_{n}\widetilde{H}^{\prime\prime\prime})$ are weak lc models of $(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$ and $(\widetilde{X},\widetilde{\Delta}+\lambda_{n}\widetilde{H})$ over $Y$ around $W$, respectively.
@@
By (i) and (iv) in Step 3 and the definition of weak lc models, $(X_{n},\Delta_{n}+\lambda^{\prime}H_{n})$ and $(X_{n},\Delta_{n}+\lambda_{n}H_{n})$ are also weak lc models of $(\widetilde{X},\widetilde{\Delta}+\lambda^{\prime}\widetilde{H})$ and $(\widetilde{X},\widetilde{\Delta}+\lambda_{n}\widetilde{H})$ over $Y$ around $W$, respectively. Let $h\colon V^{\prime}\to X_{n}$ and $\widetilde{h}\colon V^{\prime}\to\widetilde{X}^{\prime\prime\prime}$ be a common resolution of the induced bimeromorphic map $X_{n}\dashrightarrow\widetilde{X}^{\prime\prime\prime}$. By Lemma 3.3 and shrinking $Y$ around $W$, we have
$\begin{split}h^{*}(K_{X_{n}}+\Delta_{n}+\lambda^{\prime}H_{n})&=\widetilde{h}^{*}(K_{\widetilde{X}^{\prime\prime\prime}}+\widetilde{\Delta}^{\prime\prime\prime}+\lambda^{\prime}\widetilde{H}^{\prime\prime\prime}),\quad{\rm and}\\ h^{*}(K_{X_{n}}+\Delta_{n}+\lambda_{n}H_{n})&=\widetilde{h}^{*}(K_{\widetilde{X}^{\prime\prime\prime}}+\widetilde{\Delta}^{\prime\prime\prime}+\lambda_{n}\widetilde{H}^{\prime\prime\prime}).\end{split}$
@@
$h^{*}(K_{X_{n}}+\Delta_{n})=\widetilde{h}^{*}(K_{\widetilde{X}^{\prime\prime\prime}}+\widetilde{\Delta}^{\prime\prime\prime}).$
@@ 3.6. Special termination
:=$\mathcal{I}:=\left\{\frac{m-1}{m}+\sum_{j=1}^{n}\frac{r_{j}k_{j}}{m}\;\middle|\;m\in\mathbb{Z}_{>0},\,k_{j}\in\mathbb{Z}_{\geq 0}\right\}.$
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of an effective $\mathbb{R}$-divisor $A$. Let $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ be the data of the log MMP. Let $S$ be an lc center of $(X,\Delta)$ and $(S,\Delta_{S})$ the dlt pair defined by adjunction $K_{S}+\Delta_{S}=(K_{X}+\Delta)|_{S}$. Suppose that for every $i\geq 0$, after shrinking $Y$ to $Y_{i}$ so that the $i$ steps of the log MMP from the beginning are represented by a bimeromorphic contraction $X_{0}\dashrightarrow X_{i}$, then the isomorphic locus of $X_{0}\dashrightarrow X_{i}$ intersects general points of $S$. Let $S_{i}$ be the corresponding lc center of $(X_{i},\Delta_{i})$, and let $(S_{i},\Delta_{S_{i}})$ be the dlt pair defined by adjunction. We put $A_{S_{i}}:=A_{i}|_{S_{i}}$ and $A_{S_{i+1}}:=A_{i+1}|_{S_{i+1}}$. Suppose the following condition:
@@
Then there exists $i_{0}$ satisfying the following; For every $i\geq i_{0}$, after shrinking $Y$ to $Y_{i+1}$ so that the $(i+1)$-th step of the log MMP is represented by a bimeromorphic contraction $X_{i}\dashrightarrow X_{i+1}$, then the induced meromorphic map $S_{i}\dashrightarrow S_{i+1}$ is a small bimeromorphic contraction and the strict transform of $\Delta_{S_{i}}$ (resp. $A_{S_{i}}$) to $S_{i+1}$ is $\Delta_{S_{i+1}}$ (resp. $A_{S_{i+1}}$).
By shifting the log MMP, we assume the conclusion of Theorem 3.17. By shrinking $Y$ around $W$, we may assume that the first step of the MMP is represented by a bimeromorphic contraction. Over $Y$, we have diagrams
[…]
where $f_{0}$ is the $(K_{X_{0}}+\Delta_{0})$-negative extremal contraction, $f_{S_{0}}:=f_{0}|_{S_{0}}$, $f^{\prime}_{S_{0}}:=f^{\prime}_{0}|_{S_{1}}$, and $\phi_{S_{0}}:=\phi_{0}|_{S_{0}}$. By construction, $-(K_{S_{0}}+\Delta_{S_{0}})$ and $K_{S_{1}}+\Delta_{S_{1}}$ are ample over $Z_{0}$ and $\phi_{S_{0}*}\Delta_{S_{0}}=\Delta_{S_{1}}$. By [F22a, Theorem 1.21] and shrinking $Y$ around $W$, we get a dlt blow-up $g_{0}\colon(T_{0},\Gamma_{0})\to(S_{0},\Delta_{S_{0}})$ of $(S_{0},\Delta_{S_{0}})$ such that $T_{0}$ is $\mathbb{Q}$-factorial over $W$. Let $H_{Z_{0}}$ be a Cartier divisor on $Z_{0}$ which is ample over $Y$, and let $H_{S_{0}}\geq 0$ be a $\mathbb{Q}$-Cartier divisor on $S_{0}$ such that $H_{S_{0}}\sim_{\mathbb{R}}(3\cdot{\rm dim}\,X)f_{S_{0}}^{*}H_{Z_{0}}$. We put $H_{S_{1}}=\phi_{S_{0}*}H_{S_{0}}$. By choosing $H_{S_{0}}$ appropriately, we may assume that $(S_{0},\Delta_{S_{0}}+H_{S_{0}})$ and $(S_{1},\Delta_{S_{1}}+H_{S_{1}})$ are dlt pairs. Shrinking $Y$ around $W$, we may assume that $K_{S_{1}}+\Delta_{S_{1}}+H_{S_{1}}$ is ample over $Y$, and $(S_{1},\Delta_{S_{1}}+H_{S_{1}})$ is a weak lc model of both $(S_{0},\Delta_{S_{0}}+H_{S_{0}})$ and $(T_{0},\Gamma_{0}+g_{0}^{*}H_{S_{0}})$.
@@
[…]
@@
[…]
@@
$\begin{split}\lambda_{i}:=&{\rm inf}\{t\in\mathbb{R}_{\geq 0}\,|\,\text{$K_{X_{i}}+\Delta_{i}+tA_{i}$ {\rm is nef over} $W$}\},\qquad{\rm and}\\ \mu_{i}:=&{\rm inf}\{t\in\mathbb{R}_{\geq 0}\,|\,\text{$K_{T_{k_{i}}}+\Gamma_{k_{i}}+tA_{T_{k_{i}}}$ {\rm is nef over} $W$}\}.\end{split}$
@@
is a biholomorphism. By Theorem 2.26, the $(i+1)$-th step of the $(K_{X}+\Delta)$-MMP over $Y$ around $W$ does not modify a Zariski open neighborhood of $S_{i}$ for all $i\gg i_{0}$.
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of an effective $\mathbb{R}$-divisor $A$. Let $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ be the data of the log MMP. Let $S$ be an lc center of $(X,\Delta)$ and $(S,\Delta_{S})$ the dlt pair defined by adjunction $K_{S}+\Delta_{S}=(K_{X}+\Delta)|_{S}$. Suppose that for every $i\geq 0$, after we shrink $Y$ to $Y_{i}$ so that the $(K_{X_{0}}+\Delta_{0})$-MMP $(X_{0},\Delta_{0})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{0}\dashrightarrow X_{i}$, the isomorphic locus of $X_{0}\dashrightarrow X_{i}$ intersects general points of $S$. Let $S_{i}$ be the corresponding lc center of $(X_{i},\Delta_{i})$, and let $(S_{i},\Delta_{S_{i}})$ be the dlt pair defined by adjunction. Suppose the following conditions:
@@ 3.7. Lift of MMP
By shrinking $Y$ around $W$, we may assume that the first step of the MMP is represented by a bimeromorphic contraction. Over $Y$, we have a diagram
[…]
@@
[…]
@@
[…]
@@ 4.1. Definitions and basic properties
$m_{P}(D):={\rm max}\set{n\in\mathbb{Z}_{\geq 0}}{\text{$\pi_{*}\mathcal{O}_{X}(\lfloor D\rfloor-nP)\hookrightarrow\pi_{*}\mathcal{O}_{X}(\lfloor D\rfloor)$ is an isomorphism}}.$
@@
Pick $G\geq 0$ such that $D-G$ is $\pi$-ample. For any $E\in|D/Y|_{\mathbb{R}}$ and $\epsilon\in\mathbb{Q}_{>0}$, the divisor $D-E+\epsilon(D-G)$ is $\pi$-ample, and therefore we have
@@
In Lemma 4.3, the Steinness of base space is crucial. Even in the algebraic case [LX23, Section 3], the quasi-projectivity of the base variety is crucial. For example, consider a complete but not projective variety $Y$ with the trivial Picard group (see, for example, [CLS11, Example 4.2.13]). Let $X\to Y$ be a blow-up at a smooth point, and let $E$ be the exceptional divisor. Then $-E$ is clearly big over $Y$, but there is no effective $\mathbb{R}$-divisor $E^{\prime}$ such that $E^{\prime}\sim_{\mathbb{R},\,Y}-E$ because $Y$ has the trivial Picard group. In this case, the invariant
${\rm inf}\set{{\rm coeff}_{P}(E^{\prime})}{E^{\prime}\in|-E/Y|_{\mathbb{R}}}$
@@
    For a $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$ and a positive real number $r$, we have
    $\sigma_{P}(rD;X/Y)=r\sigma_{P}(D;X/Y).$
@@
    Let $W\subset Y$ be a subset. Suppose that $P$ is a prime divisor on $X$ and $\pi(P)\cap W\neq\emptyset$. For any $\mathbb{R}$-Cartier divisor $D$ on $X$, if $D$ is a limit of movable divisors over $W$ then
    $\sigma_{P}(D;X/Y)=0.$
@@
    For two $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisors $D_{1}$ and $D_{2}$ on $X$, we have
    $\sigma_{P}(D_{1}+D_{2};X/Y)\leq\sigma_{P}(D_{1};X/Y)+\sigma_{P}(D_{2};X/Y).$
@@
    For two $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisors $D_{1}$ and $D_{2}$ on $X$, we have
    $\sigma_{P}(D_{1};X/Y)=\underset{\epsilon\to 0+}{\rm lim}\sigma_{P}(D_{1}+\epsilon D_{2};X/Y).$
@@
    Let $f\colon X^{\prime}\to X$ be a projective bimeromorphism from a normal analytic variety $X^{\prime}$. For a $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$ and an effective $f$-exceptional $\mathbb{R}$-Cartier divisor $E$ on $X^{\prime}$, we have
    $\sigma_{P}(f^{*}D+E;X^{\prime}/Y)=\sigma_{P}(D;X/Y)+{\rm ord}_{P}(E).$
@@
    Suppose that $P$ is a prime divisor on $X$. Let $U\subset Y$ be a Stein open subset such that $U\cap\pi(P)\neq\emptyset$. We put $X_{U}:=\pi^{-1}(U)$ and let $P|_{X_{U}}=\bigcup_{j}P_{j}$ be the irreducible decomposition. For any $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$, all $P_{j}$ satisfy
    $\sigma_{P}(D;X/Y)=\sigma_{P_{j}}(D|_{X_{U}};X_{U}/U).$
@@
[…]
@@
[…]

### v2

@@ 3.5. From existence of weak lc model to termination
[…]
@@
[…]
@@
[…]
@@ 3.6. Special termination
$\mathcal{I}:=\left\{\frac{m-1}{m}+\sum_{j=1}^{n}\frac{r_{j}k_{j}}{m}\;\middle|\;m\in\mathbb{Z}_{>0},\,k_{j}\in\mathbb{Z}_{\geq 0}\right\}.$
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of an effective $\mathbb{R}$-divisor $A$. Let $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ be the data of the log MMP. Let $S$ be an lc center of $(X,\Delta)$ and $(S,\Delta_{S})$ the dlt pair defined by adjunction $K_{S}+\Delta_{S}=(K_{X}+\Delta)|_{S}$. Suppose that for every $i\geq 0$, after shrinking $Y$ to $Y_{i}$ so that the $i$ steps of the log MMP from the beginning are represented by a bimeromorphic contraction $X_{0}\dashrightarrow X_{i}$ over $Y$, then the isomorphic locus of $X_{0}\dashrightarrow X_{i}$ intersects general points of $S$. Let $S_{i}$ be the corresponding lc center of $(X_{i},\Delta_{i})$, and let $(S_{i},\Delta_{S_{i}})$ be the dlt pair defined by adjunction. We put $A_{S_{i}}:=A_{i}|_{S_{i}}$ and $A_{S_{i+1}}:=A_{i+1}|_{S_{i+1}}$. Suppose the following condition:
@@
Then there exists $i_{0}$ satisfying the following; For every $i\geq i_{0}$, after shrinking $Y$ to $Y_{i+1}$ so that the $(i+1)$-th step of the log MMP is represented by a bimeromorphic contraction $X_{i}\dashrightarrow X_{i+1}$ over $Y$, then the induced meromorphic map $S_{i}\dashrightarrow S_{i+1}$ is a small bimeromorphic contraction and the strict transform of $\Delta_{S_{i}}$ (resp. $A_{S_{i}}$) to $S_{i+1}$ is $\Delta_{S_{i+1}}$ (resp. $A_{S_{i+1}}$).
By shifting the log MMP, we assume the conclusion of Theorem 3.17. By shrinking $Y$ around $W$, we may assume that the first step of the MMP is represented by a bimeromorphic contraction over $Y$. Over $Y$, we have diagrams
${(X_{0},\Delta_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{0}}$$\scriptstyle{f_{0}}$${(X_{1},\Delta_{1}),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{0}}$${Z_{0}}$
${(S_{0},\Delta_{S_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{S_{0}}}$$\scriptstyle{f_{S_{0}}}$${(S_{1},\Delta_{S_{1}}),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{S_{0}}}$${Z_{0}}$
where $f_{0}$ is the $(K_{X_{0}}+\Delta_{0})$-negative extremal contraction, $f_{S_{0}}:=f_{0}|_{S_{0}}$, $f^{\prime}_{S_{0}}:=f^{\prime}_{0}|_{S_{1}}$, and $\phi_{S_{0}}:=\phi_{0}|_{S_{0}}$. By construction, $(K_{S_{0}}+\Delta_{S_{0}})$ and $K_{S_{1}}+\Delta_{S_{1}}$ are ample over $Z_{0}$ and $\phi_{S_{0}*}\Delta_{S_{0}}=\Delta_{S_{1}}$. By [F22a, Theorem 1.21] and shrinking $Y$ around $W$, we get a dlt blow-up $g_{0}\colon(T_{0},\Gamma_{0})\to(S_{0},\Delta_{S_{0}})$ of $(S_{0},\Delta_{S_{0}})$ such that $T_{0}$ is $\mathbb{Q}$-factorial over $W$. Let $H_{Z_{0}}$ be a Cartier divisor on $Z_{0}$ which is ample over $Y$, and let $H_{S_{0}}\geq 0$ be a $\mathbb{Q}$-Cartier divisor on $S_{0}$ such that $H_{S_{0}}\sim_{\mathbb{R}}(3\cdot{\rm dim}\,X)f_{S_{0}}^{*}H_{Z_{0}}$. We put $H_{S_{1}}=\phi_{S_{0}*}H_{S_{0}}$. By choosing $H_{S_{0}}$ appropriately, we may assume that $(S_{0},\Delta_{S_{0}}+H_{S_{0}})$ and $(S_{1},\Delta_{S_{1}}+H_{S_{1}})$ are dlt pairs. Shrinking $Y$ around $W$, we may assume that $K_{S_{1}}+\Delta_{S_{1}}+H_{S_{1}}$ is ample over $Y$, and $(S_{1},\Delta_{S_{1}}+H_{S_{1}})$ is a weak lc model of both $(S_{0},\Delta_{S_{0}}+H_{S_{0}})$ and $(T_{0},\Gamma_{0}+g_{0}^{*}H_{S_{0}})$.
@@
${(T_{0},\Gamma_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$${(T_{k_{1}},\Gamma_{k_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{1}}$${(S_{0},\Delta_{S_{0}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{S_{0}}}$$\scriptstyle{f_{S_{0}}}$${(S_{1},\Delta_{S_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{S_{0}}}$${Z_{0},}$
@@
${(T_{0},\Gamma_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$${(T_{k_{1}},\Gamma_{k_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(T_{k_{i}},\Gamma_{k_{i}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{i}}\textstyle{\cdots}$${(S_{0},\Delta_{S_{0}})}$${(S_{1},\Delta_{S_{1}})}\textstyle{\cdots}$${(S_{i},\Delta_{S_{i}})}\textstyle{\cdots}$
@@
[…]
@@
is a biholomorphism over $Y$. By Theorem 2.26, the $(i+1)$-th step of the $(K_{X}+\Delta)$-MMP over $Y$ around $W$ does not modify a Zariski open neighborhood of $S_{i}$ for all $i\gg i_{0}$.
@@
be a sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of an effective $\mathbb{R}$-divisor $A$. Let $\{Y_{i}\}_{i\geq 0}$ and $\{\phi_{i}\colon X_{i}\dashrightarrow X^{\prime}_{i}\}_{i\geq 0}$ be the data of the log MMP. Let $S$ be an lc center of $(X,\Delta)$ and $(S,\Delta_{S})$ the dlt pair defined by adjunction $K_{S}+\Delta_{S}=(K_{X}+\Delta)|_{S}$. Suppose that for every $i\geq 0$, after we shrink $Y$ to $Y_{i}$ so that the $(K_{X_{0}}+\Delta_{0})$-MMP $(X_{0},\Delta_{0})\dashrightarrow(X_{i},\Delta_{i})$ is represented by a bimeromorphic contraction $X_{0}\dashrightarrow X_{i}$ over $Y$, the isomorphic locus of $X_{0}\dashrightarrow X_{i}$ intersects general points of $S$. Let $S_{i}$ be the corresponding lc center of $(X_{i},\Delta_{i})$, and let $(S_{i},\Delta_{S_{i}})$ be the dlt pair defined by adjunction. Suppose the following conditions:
@@ 3.7. Lift of MMP
By shrinking $Y$ around $W$, we may assume that the first step of the MMP is represented by a bimeromorphic contraction over $Y$. Over $Y$, we have a diagram
${(X_{0},\Delta_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{0}}$$\scriptstyle{f_{0}}$${(X_{1},\Delta_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{0}}$${Z_{0}}$
@@
${(\widetilde{X}_{0},\widetilde{\Delta}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$${(\widetilde{X}_{k_{1}},\widetilde{\Delta}_{k_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{1}}$${(X_{0},\Delta_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{0}}$$\scriptstyle{f_{0}}$${(X_{1},\Delta_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f^{\prime}_{0}}$${Z_{0},}$
@@
${(\widetilde{X}_{0},\widetilde{\Delta}_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{0}}$${(\widetilde{X}_{k_{1}},\widetilde{\Delta}_{k_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{1}}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\widetilde{X}_{k_{i}},\widetilde{\Delta}_{k_{i}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{i}}\textstyle{\cdots}$${(X_{0},\Delta_{0})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{1},\Delta_{1})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(X_{i},\Delta_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots}$
@@ 4.1. Definitions and basic properties
[…]
@@
Pick $G\geq 0$ such that DGD-G is $\pi$-ample. For any $E\in|D/Y|_{\mathbb{R}}$ and $\epsilon\in\mathbb{Q}_{>0}$, the divisor $D-E+\epsilon(D-G)$ is $\pi$-ample, and therefore we have
@@
In Lemma 4.3, the Steinness of base space is crucial. Even in the algebraic case [LX23, Section 3], the quasi-projectivity of the base variety is crucial. For example, consider a complete but not projective variety $Y$ with the trivial Picard group (see, for example, [CLS11, Example 4.2.13]). Let $X\to Y$ be a blow-up at a smooth point, and let $E$ be the exceptional divisor. Then E-E is clearly big over $Y$, but there is no effective $\mathbb{R}$-divisor $E^{\prime}$ such that $E^{\prime}\sim_{\mathbb{R},\,Y}-E$ because $Y$ has the trivial Picard group. In this case, the invariant
[…]
@@
    For a $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$ and a positive real number $r$, we have σ_P(rD;X/Y)=rσ_P(D;X/Y).
@@
    Let $W\subset Y$ be a subset. Suppose that $P$ is a prime divisor on $X$ and $\pi(P)\cap W\neq\emptyset$. For any $\mathbb{R}$-Cartier divisor $D$ on $X$, if $D$ is a limit of movable divisors over $W$ then σ_P(D;X/Y)=0.
@@
    For two $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisors $D_{1}$ and $D_{2}$ on $X$, we have σ_P(D_1+D_2;X/Y) σ_P(D_1;X/Y)+σ_P(D_2;X/Y).
@@
    For two $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisors $D_{1}$ and $D_{2}$ on $X$, we have σ_P(D_1;X/Y) = ϵ0+limσ_P(D_1+ϵD_2;X/Y).
@@
    Let $f\colon X^{\prime}\to X$ be a projective bimeromorphism from a normal analytic variety $X^{\prime}$. For a $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$ and an effective $f$-exceptional $\mathbb{R}$-Cartier divisor $E$ on $X^{\prime}$, we have σ_P(f^*D+E;X’/Y)=σ_P(D;X/Y)+ord_P(E).
@@
    Suppose that $P$ is a prime divisor on $X$. Let $U\subset Y$ be a Stein open subset such that $U\cap\pi(P)\neq\emptyset$. We put $X_{U}:=\pi^{-1}(U)$ and let $P|_{X_{U}}=\bigcup_{j}P_{j}$ be the irreducible decomposition. For any $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor $D$ on $X$, all $P_{j}$ satisfy σ_P(D;X/Y)=σ_P_j(D—_X_U;X_U/U).
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$. Let $W\subset Y$ be a compact subset. Fix a resolution $\widetilde{X}\to X$ of $X$. Let $D$ be a $\pi$-pseudo-effective $\mathbb{R}$-Cartier divisor on $X$ and let $l$ be (possibly infinite) the number of prime divisors $P$ on $X$ satisfying $\pi(P)\cap W\neq\emptyset$ and $\sigma_{P}(D;X/Y)>0$. Then the inequality $l\leq\rho(\widetilde{X}/Y;W)$ holds. In particular, if $W\cap Z$ has only finitely many connected components for any analytic subset $Z\subset Y$ which is defined over an open neighborhood of $W$ (see also the condition $({\rm P}4)$ in [F22a]), then $l<\infty$.
@@
  • If $W$ is compact and $W\cap Z$ has only finitely many connected components for any analytic subset $Z$ which is defined over an open neighborhood of $W$ (see also the condition $({\rm P}4)$ in [F22a]), then $N_{\sigma}(D;X/Y,W)$ has only finitely many components.